a merchant sold 7 5/8 meters of cloths to a customer if it was cut from a bolt that contianed 18 2/3 meters
how many meters were left?

Answer by stanbon(75887)

a merchant sold 7 5/8 meters of cloths to a customer if it was cut from a bolt that contianed 18 2/3 meters
how many meters were left?
------------------
18 2/3 - 7 5/8
= 56/3 - 61/8
=(8*56 - 61*3)/24
= 265/24
= 11.04166666666... meters left
Cheers,
Stan H.